SkinnyMinnie20:

Use the net to find the lateral area of the cylinder. Give your answer in terms of n. 42π ft2 7π ft2 84π ft2 252π ft2

SkinnyMinnie20:

1 attachment
Vocaloid:

lateral area of a cylinder = the area of the rectangular part of the net so height * radius * 2 * pi = your answer

SkinnyMinnie20:

131.88

Vocaloid:

you don't have to convert to decimals 6*2*7 = 84 so your answer is 84 pi

SkinnyMinnie20:

But said use pi

Vocaloid:

I did use pi, I just didn't convert it to decimals (if you look at the answer choices they use the exact pi, not decimal pi)

SkinnyMinnie20:

oh okay

Vocaloid:

height * radius * 2 * pi = 6 * 7 * 2 * pi = 84pi = your answer